# SCAR: Scheduling Multi-Model AI Workloads on Heterogeneous Multi-Chiplet Module Accelerators

### v1

@@ V-A Experimental Settings
V-B Search Space Exploration Analysis
We compare the $3\times 3$ MCM brute-force search experiments for the heterogeneous and baseline configurations across the different optimization targets for the datacenter and AR/VR usage scenarios. All evaluations are normalized by the standalone NVDLA baseline. Though we performed experiments for all use-cases listed in Table II, we focues our analysis on a subset of experiments in Figures 7 and 8 due to space limitations: Scenarios 3 and 4 from the datacenter use cases, and the AR Gaming, Outdoors, and VR gaming scenarios from the AR/VR use-cases.
Pipelining Speedups. We observe that the combination of pipelining and chiplet heterogeneity lead to Pareto-optimal operating points that offer performance improvement opportunities. In Figure 7, we see that pipelining individually can offer speedups over standalone baselines. For example, Scenario 3 (Latency Search) – top-left most sub-Figure – shows Simba (NVDLA) realizing configurations achieving up to 4$\times$ speedup over the standalone NVDLA. This results from potential multiple chiplet assignments per window to each model, speeding up compute-intensive layers of GPT-L,BERT-L, and the 32-batch ResNet-50 models.
Heterogeneous Integration Synergy. As the density of multi-model workloads increase (Scenario 4 from Figure 7), we observe the effectiveness of homogeneous pipelining drops due to the increased competition for chiplet resources from heavier workloads. Heterogeneous MCM solutions become viable in such cases as they add another dimension for boosting performance through heterogeneous pipelining schedules, which compensate for the rising complexity through considering the varying affinities of diverse model layers, improving both latency and energy efficiency as seen in their respective search experiments. This benefits of heterogeneous pipelining also hold for the AR/VR scenarios as seen by up to 1.25×1.25$\times$1.25 × execution speedups in Figure 8.
Model Suite Diversity. The degree of diversity within the multi-model use-case also influences the overall performance improvement. In Scenario 3 (Figure 7 top), GPT-L and BERT-L were dominant transformer-based workloads with strong affinities towards the NVDLA dataflow style, contributing to the homogeneous Simba (NVDLA) solutions dominating the Pareto frontier. Scenario 4 (Figure 7 bot) and the AR/VR (Figure 8), more diversification of the layer workloads set lead more heterogeneous pipelining solutions to dominate the Pareto frontier.
Target Optimization. As illustrated in Figure 7, the dominance of MCM configurations is also dependent on the target optimization metric. For instance in Figure 7 Scenario 4, the Standalone NVDLA baseline is the most energy-efficient solution as it does not incur NoP data movement costs from pipelining. However for Energy Search experiment, Het-Sides configuration identifies scheduling solutions that become the most energy efficient ones by leveraging heterogeneity to overcome extra NoP costs.
Figure 9: Comparing latency, energy, and EDP evaluations for the top-scoring candidates from every search experiment with different optimization criteria across every scenario in Table II. Values normalized by NVDLA standalone.
Figure 10: EDP and Latency comparison for the AR/VR usage scenario on various MCM templates for the EDP search experiment.
V-C Top Performing Schedules Comparison
We compare the top-performing scheduling configurations for each MCM configuration across the various scenarios.
NoP and Inter-chiplet Pipelining. We take barplot A1 in Figure 9 as an example and and zoom in on the Simba (NVDLA) configuration (2nd category) in each scenario. In Scenarios 1-3, the small number of models and limited diversity lead the benefits of inter-chiplet pipelining to outweigh the added NoP costs, enhancing throughput and achieving latency speedups over the standalone NVDLA baseline reaching 1.4×1.4$\times$1.4 ×, 1.4×1.4$\times$1.4 ×, $\times3$, respectively. Scenarios 4 and 5 sustain larger NoP overheads due to the increased traffic contention from more workloads, causing Simba (NVDLA) 1.05×1.05$\times$1.05 × slowdown compared to the baseline in both scenarios.
Heterogeneity Pattern Choices. Across the heterogeneous MCM scheduling options in Figure 9 (Het-CB and Het-Sides), we notice that in the majority of cases, Het-Sides outperforms Het-CB. The reason being is that Het-Sides presents workloads with inter-chiplet pipelining options that can either be homogeneous or heterogeneous based on the chiplets heterogeneous arrangement . This is especially beneficial in cases where are sequences of layers that can benefit from pipelining while sharing the same dataflow affinities, unlike Het-CB which can only offer the heterogeneous pipelining option.
Scenario and Optimization Target. In all matching criteria plots (A1, B2, and C3), Het-Sides configuration at the most compute-intensive and diverse scenarios 4 and 5 consistently outperforms all baselines. For example, Scenario 4 EDP in barplot C3 is reduced by factors of 2.3$\times$ and 2.6$\times$ compared to Simba (NVDLA) and Simba (Shi-diannao), respectively, while being 9.25$\times$ less than the standalone NVDLA. We also show in Figure 10 the matching EDP barplot for the AR/VR experiments. We observe that Het-Sides option remains the most efficient option compared to the Simba baselines, achieving on average $5.2\%$ improvement over the standalone NVDLA.
Figure 11: Illustration of the top-scoring partitioning strategy for the scenario 4 experiments given the heterogeneous-sides configuration. Each window showcases the top-performing schedule within, and the mapping of models onto chiplets. Time boundaries between successive windows are computed over 500 MHz.
Het-Sides Top Scheduling Strategy. In Figure 11, we illustrate the overall scheduling strategy for the top-scoring Het-Sides solution from the EDP search in Scenario 4. The Figure depicts the per-window inner schedules and the progression of accumulative latency for processing the workloads packed into each window. The distinguishing feat from this top scheduling strategy is the non-uniformity of time windows resulting from the greedy-packing heuristic, where smaller workloads (ResNet-50) are assigned to the earlier windows at the expense of larger workloads (e.g., from BERT-L) being delayed to subsequent windows. This facilitates (i) optimizing the schedules of smaller workloads at a finer level of granularity; (ii) avoiding starvation of smaller workloads. Starting from window 2, GPT-L and BERT workloads dominate the schedule, having their segmentation and mapping strategies optimized to minimize the experienced EDP in each window. In Table III, we breakdown how the latency of each window is estimated alongside their assigned number of layers.
TABLE III: End-to-end latency breakdown in seconds for the top partitioning candidate in Figure 11. ‘Ideal‘ indicates individual model latencies unconditioned on window times.
@@ V-C Top Performing Schedules Comparison
V-D Ablation on Windowing and Chiplets Scaling
@@
Ablation Study on Time Partitioning. Using Scenario 4 and Het-Sides EDP Search experiments, we study how performance changes when varying $n_{splits}$ and repeating the experiment in Figure 12. We observe the rate of EDP improvements starts to plateau after $n_{splits}$=4 similar to our analysis in Figure 4, where EDP of the top candidate is reduced from $n_{splits}$=4 to $n_{splits}$=5 by a factor of 1.04$\times$.
@@
Ablation on Greedy Packing Algorithm. Using Scenario 4 and Het-Sides configuration, We tested the efficacy of our first-fit greedy layer packing algorithm against a uniform packing baseline, which uniformly distributes layers from each model across time windows in a uniform fashion. The Greedy layer packing algorithm was superior, improving execution speedups and energy efficiency by 21.8% and 8.6%, respectively.
Scalability. We assess the scalability of our search framework using the full 6$\times$6 Simba MCM system, where we implement an evolutionary algorithm for the SEG considering the rising problem complexity from the inclusion of more chiplets. We perform the search for our default experimental settings and Scenario 4 given $n_{splits}\in\{2,3\}$. We employ a Heterog-Cross pattern and compare it against the Simba baselines in Figure 13. We find that the Heterog-Cross top-scoring schedule outperforms those from the Simba baselines in all cases across all metrics. At $n_{splits}$=3, Heterog-Cross leads to 2.3$\times$ and 1.9 $\times$ reduction in EDP; 2.1$\times$ and 1.8$\times$ reduction in latency; over Simba (Shi-diannao) and Simba (NVDLA) and, respectively.
Figure 12: Sweep of $n_{splits}$ for the $3\times 3$ Heterog-Sides MCM configuration under EDP search showing EDP and latency results for each case’s top choice.
Figure 13: Comparing EDP and Latency at $n_{splits}\in\{2,3\}$ for the $6\times 6$ MCM for target EDP optimizaiton and an evolutionary search algorithm.
VI Related Works
Scheduler for Accelerators. Table IV compares our work against prior scheduling works. As shown, the related works can be categorized into two groups: one which has considered aspects of inter-layer pipelining and chiplet-based systems [48, 10, 52, 4, 5], while the other category of works focused on multi-model workloads on heterogeneous platforms [26, 22, 12, 30, 23]. Only this work addressed MCM, multi-model workloads, inter-layer pipelining, and heterogeneous dataflow.
Multi-chiplet Modules. Several works have proposed to address the performance scalability challenge for high-performance computing and DNN acceleration via MCM integration [48, 18, 40, 52, 1]. Simba [48] is one notable workload which pioneered a scalable deep learning inference accelerator employing MCM integration leveraging non-uniform work partitioning, communication-aware data placement, and cross-layer pipelining.
Intra- and Inter-layer Parallelism. Numerous works have explored intra-layer parallelism to maximize performance efficiency and resource utilization by partitioning DNN layers into smaller parallelizable tiles [41, 15, 17, 31, 58, 57, 16]. Other works have studied the inter-layer scheduling space to compensate for workloads characterized by low degrees of parallelism [10, 21, 33, 60, 4].
TABLE IV: Comparison against prior related scheduling works.
@@ VI Related Works
VII Conclusion
In this work, we explored the scheduling space of a new class of MCM accelerator architecture, heterogeneous MCM AI accelerator, targeting multi-model AI workloads. We identify that the scheduling problem is intractably large but multi-level problem formulation and heuristics we proposed are effective for the extended scheduling problem. The results also show that heterogeneous MCM accelerator is beneficial multi-model workloads, which motivates further exploration.

### v2

@@ V-C AR/VR multi-model scheduling results
TABLE V: EDP search AR/VR results normalized by standalone NVDLA
Strgy EDP Search
Relative Latency Relative EDP
Sc6 Sc7 Sc8 Sc9 Sc10 Sc6 Sc7 Sc8 Sc9 Sc10
Stand.(Shi) 1.7 1.7 0.96 0.36 1.24 2.32 2.26 1.22 0.45 1.71
Stand.(NVD) 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0
Simba(Shi) 1.72 1.75 0.97 0.36 1.28 2.36 2.33 1.24 0.47 1.78
Simba(NVD) 1.01 1.0 0.99 1.07 1.0 1.01 1.01 1.0 1.08 1.02
Het-CB 1.01 1.03 0.78 0.4 0.77 1.09 1.09 0.83 0.43 0.82
Het-Sides 1.0 0.99 0.78 0.41 0.75 1.05 1.03 0.83 0.45 0.79
TABLE VI: End-to-end latency breakdown in seconds for the top partitioning candidate in Figure 9.
@@
Figure 12: Results for the EDP Search for Scenarios 3 and 4 on the Triangular NoP topologies from Figure 6 (Normalized by standalone NVDLA).
V-D Scaling to 6$\times$6 MCM system
We assess how SCAR scheduling scales with the hardware using the 6$\times$6 full Simba MCM system with its templates from Figure 6. We implement an evolutionary algorithm for the SEG engine to scale with the rising problem size, and show the results for the EDP search on Scenario 4 at $n_{splits}$=$2$ and $n_{splits}$=$3$. We choose the heterogeneous cross (Het-Cross) as our heterogeneous template following the insights from the 3$\times$3 experiment regarding choosing heterogeneity patterns that enable both homogeneous and heterogeneous pipelining capabilities. As shown in Figure 13, the evolutionary search has led to identifying configurations for Het-Cross that achieve 2.3$\times$ and 1.9$\times$ reduction in EDP; 2.1$\times$ and 1.8$\times$ reduction in latency over Simba (Shi) and Simba (NVD), respectively.
Figure 13: Pareto Plot and comparison of top performing models for the EDP search on the $6\times 6$ MCM at $n_{splits}$=2 (top) and $n_{splits}$=3 (bot).
V-E Ablation Studies
@@
Ablation Study on Time Partitioning. Using Scenario 4 and Het-Sides strategy, we study how performance changes when varying $n_{splits}$ from 1 to 5 and repeating the EDP search experiment. Prior to $n_{splits}$=4, the average rate of reduction in EDP was 1.25$\times$. The rate of improvement drops to only 1.04$\times$ between $n_{splits}$=4 to 5, indicating diminishing returns.
Ablation Study on other NoP Topology. SCAR can generalize to other NoP topologies as it relies on adjacency matrix connectivity. We test this by performing the EDP search for Scenarios 3 and 4 using the triangular NoP topologies in Figure 6. As shown in Figure 12, similar performance patterns are exhibited compared to the Scenario results on the 3$\times$3 Mesh, albeit with varying relative gains. For instance, though Het-T outperforms both Simba-T strategies in Scenario 4 by 2.5$\times$ and 1.67$\times$ over Simba-T (Shi) and Simba-T (NVD), respectively, it is second best compared to Simba-T (NVD) in Scenario 3 by a factor of 2.5$\times$ (compared to 1.47$\times$ from Figure 7-C3) due to the increased competition for resources.
Ablation Study on rule-based PROV. We repeated the EDP search for all strategies across Scenarios 3-5 on the $3\times 3$ template using an exhaustive search over the $N_{i}$ values. Though the results improved further with the added search complexity, the insights remained the same. For Scenarios 4-5, Het-Sides remained superior, achieving respective EDP reductions of 38.3% and 29.9% from Standalone NVDLA; 59.5% and 57.6% from Simba (Shi); 33.0% and 28.3% from Simba (NVD). For Scenario 3, Simba (NVD) remained superior, achieving 79.7% EDP reduction Het-Sides. These evaluations follow the trends from our rule-based results in Table IV.
@@
Ablation on Greedy Packing Algorithm. Using Scenario 4 and Het-Sides, we test the efficacy of our first-fit greedy layer packing algorithm against a uniform packing baseline, distributing layers uniformly across time windows. Ours achieved 21.8% speedup and 8.6% energy reduction.
V-F Summary of Results and Main Insights
We summarize our main insights and findings as follows.
  • Heterogeneous MCM patterns improve performance for heavy and diverse multi-model workloads (scenarios 4-5).
  • Homogeneous MCM patterns are more suited for small multi-model workloads (scenarios 1-3).
  • Heterogeneous MCM patterns with diverse pipelining options (Het-Sides) are superior to heterogeneous patterns with homogeneous pipelining options (Het-CB).
  • The target optimization objective is crucial in identifying the best integration strategy. In EDP search scenario 4, Het-sides outperformed all other strategies on EDP, but not on pure energy consumption.
  • Topology and number of resources affect the extent of performance improvement for heterogeneous strategies.
Our findings show that understanding multi-model workload characteristics and usage scenarios is crucial for identifying the best MCM integration strategy for a target objective.
VI Discussion and Limitations
Multi-model optimization targets. We experimented with different optimization targets (latency, EDP, energy) for our scenarios, and showed that the top performing strategy can change based on the target objective. As multi-model workloads evolve, it may be desirable to assign separate optimization targets for different models within a scenario (EDP v. lat). One practical way to achieve this in our framework is by adding a constraint in our EDP search, invalidating schedules that have certain models violate a latency constraint (i.e., the EDP search becomes lower bounded by the latency search).
Heterogeneous chiplets technology. Heterogeneous chiplet integration has become a viable, cost-effective approach to design state-of-the-art AI systems. Nvidia’s world-class superchips are a successful example of heterogeneous on-package integration (e.g., Grace-Blackwell (1 CPU + 2 GPUs) [14]). The success of these systems and others (AMD’s MI300X [1]) is testament to the hardware manufacturers’ investment in chiplets technology, where through advanced manufacturing processes and heterogeneous integration capabilities, the development of MCM AI accelerators (like Nvidia’s Simba [64]) becomes more accessible, allowing chiplet modifications/replacement in MCM hardware at lower costs without requiring a complete overhaul of the entire package.
Scheduler Software Integration. SCAR can be integrated on top of existing compiler infrastructure. The advanced scheduling techniques supported by the scheduler (dynamic chiplet regrouping, inter-chiplet pipelining) represent high-level abstractions of the computational graphs that can be transformed through standard compiler software (e.g., MLIR [39]) to representations suited for the underlying hardware. For example, dynamic chiplets regrouping is correspondent to graph partitioning, where a model’s computational graph is divided into smaller subgraphs, each associated with the set of computing nodes assigned during the corresponding time window. The subgraphs can then be transformed to lower representations covering the details of buffer management, die-to-die communication, memory R/W requests, I/O, all the way to the transformations covering the dataflow features (loop reordering, spatial unrolling) for the specialized accelerators.
VII Related Works
TABLE VII: Comparison against prior related scheduling works.
@@ VII Related Works
Scheduler for Accelerators. Table VII compares our work against prior scheduling works. As shown, the related works can be categorized into two groups: one which has considered aspects of inter-layer pipelining and chiplet-based systems [64, 19, 68, 8, 9], and another that focused on multi-model workloads on heterogeneous platforms [37, 32, 21, 42, 33]. Only this work addressed MCM, multi-model workloads, inter-layer pipelining, and heterogeneous dataflow.
Multi-chiplet Modules. Several works proposed to address the scalability challenge for DNN acceleration via MCM integration [64, 28, 55, 68, 3]. Most notably, Simba [64] pioneered a scalable deep learning MCM inference accelerator leveraging non-uniform work partitioning, communication-aware data placement, and cross-layer pipelining.
Intra- and Inter-layer Parallelism. Prior works explored intra-layer parallelism to maximize DNN performance efficiency by partitioning DNN layers into smaller, parallelizable tiles [57, 25, 27, 43, 74, 73, 26]. Other works studied the inter-layer scheduling space to compensate for workloads with low degrees of parallelism [19, 31, 45, 78, 8, 53, 6].
VIII Conclusion
In this work, we explored the scheduling space of a new class of MCM accelerator architecture, heterogeneous MCM AI accelerator, targeting multi-model AI workloads. We identify that the scheduling problem is intractably large but multi-level problem formulation and heuristics we proposed are effective for the large-scale scheduling problem. The results also show that heterogeneous MCM accelerator is beneficial for multi-model workloads, which motivates further exploration.